OpenStudy (anonymous):

Use the graph below to answer the question that follows: graph of curve that passes through the following points, 0, 5; pi over 2, 2; pi, negative 1; 3 pi over 2, 2; 2 pi, 5. What is the rate of change between the interval x = π and x = three pi over two? pi over six six over pi 4 over pi pi over 4 @SolomonZelman Lplz plz plz help me asap

OpenStudy (anonymous):

I NEED HELP PLZZ

OpenStudy (anonymous):

got a picture?

OpenStudy (anonymous):

The highest point goes from (0, 5) down to (pi/2,2) back down to (pi,-1) up to (3pi/2,2) and back up to( 2pi,5)

OpenStudy (anonymous):

rate of change would be \[\frac{2+1}{\frac{3\pi}{2}-\pi}\] which is none of your choices

OpenStudy (anonymous):

oh maybe it is

OpenStudy (anonymous):

\[\frac{3}{\frac{\pi}{2}}=\frac{6}{\pi}\] is that an option?

OpenStudy (anonymous):

yeah it is go with B

OpenStudy (anonymous):

thankss!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Simplify cos2 Θ(1 + tan2 Θ). cos2 Θ + sin2 Θ cos2 Θ 1 −1 @satellite73

OpenStudy (anonymous):

\[\cos^2(\theta)(1+\tan^2(\theta))\]

OpenStudy (anonymous):

thank you sooo much!

OpenStudy (anonymous):

Simplify (1 − sin x)(1 + sin x). 1 cos2 x sin2 x tan2 x

OpenStudy (anonymous):

use the identity \[1+\tan^2(\theta))=\sec^2(\theta)\]

OpenStudy (anonymous):

hold on got the question marks fo death again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

if you multiply you get \[1-\sin^2(\theta)\] as a first step yeah B

OpenStudy (anonymous):

Compare the functions shown below: f(x) = 3 cos 2x + 4 g(x) cosine graph with points at 0, 3 and pi over 2, 0 and pi, negative 3 and 3 pi over 2, 0 and 2 pi, 3 h(x) x y -6 -11 -5 -6 -4 -3 -3 -2 -2 -3 -1 -6 0 -11 Which function has the greatest maximum y-value? f(x) g(x) h(x) All three functions have the same maximum y-value.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

maximum we want right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hard for me to tell without the picture, but the maximum of cosine is 1, so the maximum of \(3\cos(2x)+4\) is \(3+4=7\)1

OpenStudy (anonymous):

look at the graph of \(g\) and find the top most point let me know if it is more or less than 7

OpenStudy (anonymous):

its 3 less

OpenStudy (anonymous):

ok then pick that on \(f\) because the max of \(h\) looks like \(-2\)

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Compare the functions shown below: f(x) = 4x + 5 g(x) tangent function with y intercept at 0, 2 h(x) = 3 sin(2x + π) - 2 Which function has the greatest y-intercept? f(x) g(x) h(x) All three functions have the same y-intercept.

OpenStudy (anonymous):

Compare the functions shown below: f(x) polynomial graph with x intercepts at negative 3, negative 2, 1, 2, y intercept at negative 12 g(x) trig graph with points at 0, 3 and pi over 2, 0 and pi, negative 3 and 3 pi over 2, 0 and 2 pi, 3 Over the interval x = 0 to x = 2π h(x) = 2x3 − 2x2 − 3x + 2 Which function has the most x-intercepts? f(x) g(x) h(x) All three functions have the same number of x-intercepts.

OpenStudy (anonymous):

Compare the functions shown below: f(x) x y -3 -27 -2 -8 -1 -1 0 0 1 1 2 8 3 27 4 64 g(x) linear graph with y intercept of negative 1 and x intercept of negative 2 h(x) = (x + 4)2 + 2 What is the correct order of the functions from least to greatest according to the average rate of change on the interval from x = 0 to x = 4? f(x), g(x), h(x) g(x), f(x), h(x) h(x), g(x), f(x) g(x), h(x), f(x)

OpenStudy (anonymous):

Compare the functions shown below: f(x) = 2 sin (3x + π) − 2 g(x) = (x − 3)2 − 1 h(x) x y -2 3 -1 -2 0 -5 1 -6 2 -5 3 -2 4 3 Which function has the smallest minimum y-value? f(x) g(x) h(x) Both g(x) and h(x) have the same minimum y-value.

OpenStudy (anonymous):

this is not working for me, too many question marks

OpenStudy (anonymous):

refresh the page and you'll be able to see them without any question marks

OpenStudy (anonymous):

@ganeshie8 can u plz help me?

ganeshie8 (ganeshie8):

looks like you're given 3 functions and asked to find which function has minimum y-value

OpenStudy (anonymous):

no on a nother question

ganeshie8 (ganeshie8):

Okay can you post it as new question.. ? :)

OpenStudy (anonymous):

alright :)